OpenStudy (anonymous):

square root of a3b7z3

OpenStudy (anonymous):

\[\sqrt{a^{3}b^{7}z^{3}} = \sqrt{a \times a \times a \times b \times b \times b \times b \times b\times b \times b \times z \times z \times z} = \]\[\sqrt{\cancel{a} \times \cancel{a} \times a \times \cancel{b} \times \cancel{b} \times \cancel{b} \times \cancel{b} \times \cancel{b}\times \cancel{b} \times b \times \cancel{z} \times \cancel{z} \times z} = ab^{3}z\sqrt{abz}\]